# Learning from Imperfect Demonstrations through Dynamics Evaluation

### 18 Dec 2023 v1

@@ Abstract
Standard imitation learning usually assumes that demonstrations are drawn from an optimal policy distribution. However, in real-world scenarios, every human demonstration may exhibit nearly random behavior and collecting high-quality human datasets can be quite costly. This requires imitation learning can learn from imperfect demonstrations to obtain robotic policies that align human intent. Prior work uses confidence scores to extract useful information from imperfect demonstrations, which relies on access to ground truth rewards or active human supervision. In this paper, we propose a dynamics-based method to evaluate the data confidence scores without above efforts. We develop a generalized confidence-based imitation learning framework called Confidence-based Inverse soft-Q Learning (CIQL), which can employ different optimal policy matching methods by simply changing object functions. Experimental results show that our confidence evaluation method can increase the success rate by $40.3\%$ over the original algorithm and $13.5\%$ over the simple noise filtering.
@@ I INTRODUCTION
Imitation learning (IL) attracts much attention for its potential to enable robots to complete complex tasks from human demonstrations. Traditional IL typically depends on large-scale, high-quality, and diverse human datasets to achieve satisfactory results[1, 2, 3]. In real-world scenarios, collecting such human datasets can be quite costly. When using specialized data collection system, human operators may provide noisy demonstrations due to factors such as fatigue, distraction, or lack of skill[4]. Each human demonstration has the potential to exhibit nearly random behavior. The well performing trajectories may still contain some noise, which can have a negative impact on the performance of standard IL[5]. The poorly performing or even unsuccessful trajectories may contain useful information. It is difficult for humans to provide perfect trajectories, and these imperfect demonstrations can not accurately reflect the expert’s true intent. In this paper, we aim to maximize the use of limited human imperfect demonstrations to learn policy that align human intent.
@@
Learning from imperfect demonstrations is nothing new, and its work falls into two main directions: ranking-based and confidence-based. For the ranking-based approach, it learns a reward function that satisfies the ranking of trajectories through supervised learning, and then improves the policy through reinforcement learning[6, 7, 8]. This approach depends on the availability of high-quality datasets to learn a robust reward function. For the confidence-based approach, it uses confidence scores to describe the quality of demonstrations in order to extract useful information from imperfect demonstrations while filtering noise efficiently[9, 10, 11]. The key challenge with this approach is how to get the true data confidence scores.
@@
Figure 1: Overview of Confidence-based Inverse soft-Q Learning (CIQL): (1) Use the confidence function $w$ to label the transitions $\{s,a,s^{\prime}\}$ and estimate the optimal priori probability $\alpha$ (2) Sample a batch of transitions from the expert and agent replay buffers, learn the optimal Q-function via CIQL, and improve the policy via Soft Actor-Critic (SAC). (3) Recover the reward function as obtain the optimal policy and evaluate whether it aligns human intent.
@@
It is impossible for humans to evaluate every transition in a trajectory, and a single confidence score for a segment or the whole trajectory can be quite superficial. Methods for obtaining fine-grained confidence in trajectories: (1) One assumes that experts can have a stable performance in proficient skills[3]. In practice, there may be random noise in each demonstration, and even an expert may make different decisions in the same state. (2) It is possible to train a classifier using partially labeled data to predict confidence scores for unlabeled data[10]. This two-step training method may accumulate errors, which can degrade performance. (3) It can also be evaluated using the reward function recovered by inverse reinforcement learning[11, 9]. Due to the ambiguity of the reward function, this method may obtain inaccurate confidence scores, which can also degrade performance[12]. In summary, these methods either require strict assumptions, active human supervision, or access to ground truth rewards. Besides this, the confidence scores obtained are not always accurate.
@@
To mitigate the above problem, we propose to evaluate the data confidence score using state transitions, which only requires setting the noise angle. There is an optimal threshold for the noise angle, and we find its general range experimentally. As shown in Fig.1, based on the state-of-the-art imitation learning, we integrate a generalized confidence-based imitation learning framework called Confidence-based Inverse soft-Q Learning (CIQL) to validate our evaluation method. Our main contributions are as follows:
@@
  • We propose a dynamics-based method to divide data into useful and noisy data by setting the noise angle. It utilizes state transition to determine the approach angle and calculates its confidence score from a confidence function that joints the approach angle with the noise angle.
  • Our framework can use different optimal policy matching methods by simply changing the objective function. Such as CIQL is implemented by estimating the Expert’s optimal policy distribution (CIQL-E) or the Agent’s non-optimal policy distribution (CIQL-A).
@@
  • We experimentally show that our confidence score evaluation method improves the performance of the original algorithm, and CIQL-A is more aligned to human intent.
@@ III-A Background: General Inverse RL Objective
Note that although IQ-Learn’s new objective does not directly recover the reward function during training, it can still do IRL by recovering the reward function via $r=\mathcal{T}^{\pi}Q$ if you want. This makes it possible to use different confidence evaluation methods, even those that require a reward function[9, 11, 27].
@@ III-B Confidence-based IQ-Learn (CIQL)
It can be interpreted that the expert policy distribution is a mixture distribution of an optimal policy with a ratio of $\alpha$ and non-optimal policies with a ratio of $1-\alpha$. Confidence-based Imitation Learning aims to minimize the statistical distance between $\rho_{\pi}$ and $\rho_{opt}$ to match the optimal policy. We assume that true confidence scores are available for all data. There are two methods to guide confidence-based policy learning.
CIQL-Expert: The first method is to estimate the optimal occupancy measure of the expert, $\rho_{opt}$. We can replace $\mathbb{E}_{\rho_{opt}}[\phi(\mathcal{T}^{\pi}Q)]$ with $\mathbb{E}_{\rho_{E}}[\phi(\mathcal{T}^{\pi}Q)]$ to directly match $\rho_{opt}$. Consequently, the objective of estimating the Expert optimal policy distribution (CIQL-E) using Eq.4 follows:
@@
CIQL-Agent: The second method is to estimate the non-optimal occupancy measure of the agent, $\rho_{\pi^{\prime}}$. It assumes that the policy distribution structure of an agent is the same as that of an expert, which can be expressed as a mixture of $\rho_{opt}$ and $\rho_{non}$, specifically $\rho_{\pi^{\prime}}=\alpha\rho_{\pi}+(1-\alpha)\rho_{non}$. Recalling $\rho_{E}=\alpha\rho_{opt}+(1-\alpha)\rho_{non}$, the statistical distance between $\rho_{opt}$ and $\rho_{\pi}$ can minimized by minimizing the statistical distance between $\rho_{\pi^{\prime}}$ and $\rho_{E}$. Then we can replace $\mathbb{E}_{\rho_{\pi^{\prime}}}[\mathcal{T}^{\pi}Q]$ with $\alpha\mathbb{E}_{\rho_{\pi}}[\mathcal{T}^{\pi}Q]+(1-\alpha)\mathbb{E}_{\rho_{non}}[\mathcal{T}^{\pi}Q]$. Consequently, the objective of estimating the Agent non-optimal policy distribution (CIQL-A) using Eq.5 follows:
@@
Using statistical distances $\phi$ such as KL, Hellinger, $\mathcal{X}^{2}$ divergence, etc., when sampling noisy data with a confidence score $w=0$, the first term $\mathbb{E}_{\rho_{E}}[\phi(\mathcal{T}^{\pi}Q)-(1-w)\cdot(\mathcal{T}^{\pi}Q)]\leq 0$ indicates that penalizing the noise. This means that the method not only allows the agent to learn how to behave from experts, but also avoids bad behavior from noise.
Input: Human data $\mathcal{D}_{E}$, noise angle $\theta_{a}$, ratio $\mu$
@@
2       Sample a batch of transitions $\{s,a,s^{\prime},w,t\}$ from $\mathcal{D}_{E}$ and $\mathcal{D}_{A}$, where $t$ is one-hot encoding of the data type Train the Q-function $Q_{\psi}$ using the objective from Eq.7 (CIQL-E) or Eq.8 (CIQL-A): $\psi_{t+1}\leftarrow\psi_{t}+l_{\psi}\cdot\nabla_{\psi}L(Q_{\psi})$ Improve the policy $\pi_{\varphi}$ using the objective from Eq.9 (SAC Actor Update):    $\varphi_{t+1}\leftarrow\varphi_{t}-l_{\varphi}\cdot\nabla_{\varphi}J(\pi_{\varphi})$ Policy $\pi$ interacts with $\mathcal{M}$ and add the new transitions to $\mathcal{D}_{A}$
@@
III-C Algorithm Implementation
To solve the CIQL problem, we need to sample a set of transitions $\{s,a,s^{\prime}\}$ data and compute the value of $\mathcal{T}^{\pi}Q$ according to the formula $Q(s,a)-\gamma\mathbb{E}_{s^{\prime}\sim P(s,a)}V^{\pi}(s^{\prime})$, as in Eq.2. The soft state value function $V^{\pi}(s^{\prime})$ is computed using the data $\{s^{\prime},a^{\prime}\}$ via its formula on the Q-function, where the next action $a^{\prime}$ is sampled from the policy $\pi$ on the next state $s^{\prime}$. In our experiments, we use the same statistical distance ($\mathcal{X}^{2}$-divergence) as in the paper[24], specifically, $\phi(x)=x-\frac{1}{4\sigma}x^{2}$ with $\sigma=0.5$. Algorithm 1 introduces our framework which consists of four main steps:
  • Calculate the confidence scores $w$ for each transition in $\mathcal{D}_{E}$ using Eq.13 and estimate the prior probability of the optimal policy $\alpha\approx\frac{1}{n}\sum_{i=1}^{n}w_{i}$.
  • For a fixed $\pi$, optimize $Q$ by maximizing $L(\pi,Q)$ from Eq.7 (CIQL-E) or 8 (CIQL-A).
  • For a fixed $Q$, improve $\pi$ by minimizing the Actor’s objective function $J(\pi)$ in Soft Actor-Critic (SAC)[26].
  • For the optimal policy $\pi^{*}$, recover the reward function under the optimal $Q^{*}$-function by applying inverse soft Bellman operator $\mathcal{T}^{\pi}$.
III-D Dynamics Evaluation
@@
Confidence Score Setting: In the reaching stage, the primary goal is for the robot to approach the target directly. This intention can be represented by the environment state reward function, as shown in Fig.2. To generate this reward function, the useful data is defined as the yellow segments in the trajectory, while the noisy data is defined as the blue segments. Our method is to utilize the approach angle $\theta_{a}$ as a classification feature, which is defined as:
$\theta_{a}=arccos\frac{{\tilde{s}_{t}o}\cdot{\tilde{s}_{t}\tilde{s}_{t+1}}}{|{\tilde{s}_{t}o}||{\tilde{s}_{t}\tilde{s}_{t+1}}|}$ (11)
where the value domain of $\theta_{a}$ is $[0,\pi]$. $\tilde{s}_{0}$ and $o$ are the initial positions of gripper and target, respectively. The data $\{\tilde{s}_{t},\tilde{s}_{t+1},o\}$ are some features in the transition $\{s_{t},a_{t},s_{t+1}\}$. For a given transition, it has a consistent value $\theta_{a}$ that can be used to calculate its confidence score $w$. Since the robot environmental dynamics are deterministic, $w(s_{t},a_{t},s_{t+1})$ is equivalent to $w(s_{t},a_{t})$. Therefore, we can categorize the data into three components based on $\theta_{a}$: optimal, non-optimal and noise.
$w(\theta_{a})=\left\{\begin{array}[]{lll}1&for\quad\theta_{a}=0,&opt\\ (0,1)&for\quad 0<\theta_{a}\leq\theta_{n},&non\ opt\\ 0&for\quad\theta_{n}<\theta_{a}\leq\pi,&noise\end{array}\right.$ (12)
where the noise angle $\theta_{n}$ is used to determine whether the data is noisy. Transitions that satisfy $\theta_{a}\leq\theta_{n}$ are considered useful, while others are considered noise. In the useful datasets, it can be further subdivided into optimal and non-optimal data. To satisfy the condition 12, we find a simple function for $w$ that significantly improves the performance of the original algorithm:
$w(\theta_{a})=\rm Sigmoid((\theta_{a}/\theta_{n}-0.5)*\mu)$ (13)
where the ratio $\mu$ satisfies $\rm Sigmoid(0.5*\mu)<\epsilon$, and $\epsilon$ is the limiting distance from the actual score to the confidence bound. In the operating stage, it is important to emphasize that it consists of few button-pressing data points. We chose to keep all the data from this stage in the demonstrations where the task can be completed successfully. Their confidence scores are set in the [1, 2] interval to emphasize that stage. Note that the confidence function $w$ relies only on dynamics information, eliminating the need for other inputs such as the ground truth rewards or active human supervision.
@@
Figure 2: The blue gradient circle is the operator’s ideal reward function (true intent) where the inner portion is high reward and the outer portion is low reward. $\tau_{1},\tau_{2},\tau_{3}$ are complete trajectories of the approaching stage. $\theta_{a}$ is the approach angle between vectors ${\tilde{s}_{t}o}$ and ${\tilde{s}_{t}\tilde{s}_{t+1}}$, and $\theta_{n}$ is the noise angle to be set.
Figure 3: Teleoperation Demonstration System. Better datasets horizons: 100-150; Worse datasets horizons: 200-400; The failure of Failed datasets is due to inability to grasp target within 500 horizon range.
@@ IV-A Experiments Setting
We develop a robotic teleoperation system using the input signaling capability of the Omega.3 haptic device to collect human demonstrations. Our system integrates seamlessly with Robosuite, which provides an IIWA robotic arm with an OSC-POSE controller. The control inputs contain a set of features including position vectors $\{x,y,z\}$, pitch, yaw, and roll angles $\{\theta_{p},\theta_{y},\theta_{r}\}$, and grasping binary signals $\{g\}$. We collect three types of human datasets at a control frequency of 20 Hz with a horizon limitation of 500: Better, Worse, and Failed. Each type consists of 30 trajectories with different horizons, as shown in Fig.3.
Our training hyperparameters are set as follows: the learning rates for both actors and critics are set to 5e-6, with a total training step of 500K. During training, we set the horizon to 200 and choose the moving average success rate of 10 trajectories as the metric for selecting the best model. During evaluation, we measure the average of the success rate for every 100 trajectories among 10 random seeds as the metric for evaluating the policy, with a set horizon of 500.
Figure 4: Effect of Noise Angle. We use the Better-Worse-Failed datasets (90) to explore the effect of noise angles on the success rates. IQ-Learn(filter) is learning on filtered noisy data, and it becomes IQ-Learn when $\theta_{a}$ equal to 180°.

### 22 Feb 2024 v2

# Aligning Human Intent from Imperfect Demonstrations

@@ Abstract
Standard imitation learning usually assumes that demonstrations are drawn from an optimal policy distribution. However, in the real world, where every human demonstration may exhibit nearly random behavior, the cost of collecting high-quality human datasets can be quite costly. This requires robots to be able to learn from imperfect demonstrations and thus acquire behavioral policy that align human intent. Prior work uses confidence scores to extract useful information from imperfect demonstrations, which relies on access to ground truth rewards or active human supervision. In this paper, we propose a dynamics-based method to obtain fine-grained confidence scores for data without the above efforts. We develop a generalized confidence-based imitation learning framework called Confidence-based Inverse soft-Q Learning (CIQL), which can employ different policy learning methods by changing object functions. Experimental results show that our confidence evaluation method can increase the success rate of the original algorithm by $40.3\%$, which is $13.5\%$ higher than the method of just filtering noise.
@@ I INTRODUCTION
Imitation learning (IL) attracts much attention for its potential to enable robots to complete complex tasks from human demonstrations. However, traditional IL typically depends on large-scale, high-quality, and diverse human datasets, which can be expensive to collect[1, 2, 3]. Real-world data collection can introduce noise into operator demonstrations, caused by fatigue, distractions, or insufficient skills [4]. Humans struggle to provide perfect trajectories, with demonstrations often exhibiting nearly random behavior. The well-performing trajectories contain noise that can negatively affect standard IL performance[5]. The poorly performing or unsuccessful trajectories may still contain useful information. These imperfect demonstrations can not accurately reflect the expert’s true intent. In this paper, we aim to maximize the use of limited imperfect human demonstrations to learn a policy aligned with human intent.
@@
Learning from imperfect demonstrations is nothing new, and its work falls into two main directions: ranking-based and confidence-based. Ranking-based methods learn a reward function through supervised learning to satisfy trajectory rankings, then improves the policy via reinforcement learning [6, 7, 8]. This approach requires high-quality datasets to learn a robust reward function. Confidence-based methods use confidence scores to describe demonstration quality, extracting useful information while efficiently filtering noise [9, 10, 11]. The challenge with this approach is obtaining true data confidence scores.
@@
Figure 1: Overview of Confidence-based Inverse soft-Q Learning (CIQL): (1) Use the confidence function $w$ to label the transitions $\{s,a,s^{\prime}\}$ and estimate the optimal priori probability $\alpha$ (2) Sample a batch of transitions to learn the optimal Q-function via CIQL, and improve the policy via Soft Actor-Critic (SAC). (3) Recover the reward function as obtain the optimal policy and evaluate whether it aligns human intent.
@@
Evaluating every transition in a trajectory is impractical for humans, and a single confidence score for a segment or the whole trajectory is insufficient. To obtain fine-grained confidence in trajectories, methods include: (1) Assuming stable expert performance[3] can be problematic due to random noise and inconsistent decisions. (2) Training a classifier with partially labeled data to predict unlabeled data confidence scores[10] can lead to accumulated errors and degraded performance. (3) Using inverse reinforcement learning to recover a reward function for evaluating data confidence scores[11, 9] may yield inaccuracies due to reward function ambiguity, also degrading performance[12]. In summary, these methods often rely on strict assumptions, active human supervision, or ground truth rewards, and may not provide accurate confidence scores.
@@
To mitigate the above problem, we introduce a method for evaluating data confidence scores using state transitions, which only requires setting the noise angle. We identify the optimal threshold range for the noise angle experimentally. As shown in Fig.1, building on state-of-the-art imitation learning, we integrate a generalized confidence-based imitation learning framework, termed Confidence-based Inverse soft-Q Learning (CIQL), to validate our evaluation approach. Our main contributions are:
@@
  • We develop a dynamics-based method to differentiate between useful and noisy data. This method leverages state transition to determine the approach angle and employs a confidence function to compute its confidence score.
  • Our framework is adaptable to various optimal policy learning methods by modifying the objective function. CIQL-E estimates the distribution of experts’ optimal policies, while CIQL-A estimates the distribution of agents’ non-optimal policies.
@@
  • Our experimental results demonstrate that our confidence score evaluation method enhances the performance of the baseline algorithm, and CIQL-A is more aligned to human intent.
@@ III-A Background: General Inverse RL Objective
Note that although IQ-Learn’s new objective does not directly recover the reward function during training, it can still do IRL by recovering the reward function via $r=\mathcal{T}^{\pi}Q$ if you want. For the optimal policy $\pi^{*}$, recover the reward function under the optimal $Q^{*}$-function by applying inverse soft Bellman operator $\mathcal{T}^{\pi}$.
This enables the use of different confidence evaluation methods, even those that require a reward function[9, 11, 27].
@@ III-B Confidence-based IQ-Learn (CIQL)
Confidence-based Imitation Learning aims to minimize the statistical distance between $\rho_{\pi}$ and $\rho_{opt}$ to match the optimal policy. There are two methods to guide confidence-based policy learning.
CIQL-Expert: The first method estimates the expert’s optimal occupancy measure using the Eq.5. The objective for estimating the expert’s optimal policy distribution (CIQL-E) is:
@@
CIQL-Agent: The second method estimates the agent’s non-optimal occupancy measure using the Eq.6. It assumes that the agent’s policy distribution structure is the same as the expert’s, given by $\rho_{\pi^{\prime}}=\alpha\rho_{\pi}+(1-\alpha)\rho_{non}$. Minimizing the statistical distance between $\rho_{\pi^{\prime}}$ and $\rho_{E}$ equates to minimizing that between $\rho_{opt}$ and $\rho_{\pi}$. The objective for estimating the agent’s non-optimal policy distribution (CIQL-A) is:
@@
when using statistical distances $\phi$ such as KL, Hellinger, $\mathcal{X}^{2}$-divergence, etc., and sampling noisy data with a confidence score $w=0$, the first term $\mathbb{E}_{\rho_{E}}[\phi(\mathcal{T}^{\pi}Q)-(1-w)\cdot(\mathcal{T}^{\pi}Q)]\leq 0$ indicates that penalizing the noise. This means that the method not only allows the agent to learn how to behave from experts, but also avoids bad behavior from noise.
In our experiments, we use the same statistical distance ($\mathcal{X}^{2}$-divergence) as in the paper[24], specifically, $\phi(x)=x-\frac{1}{4\sigma}x^{2}$ with $\sigma=0.5$. Algorithm 1 introduces our framework, where the policy improvement uses the Actor’s objective function in Soft Actor-Critic (SAC)[26].
Input: Human data $\mathcal{D}_{E}$, noise angle $\theta$, ratio $\mu$
@@
2       Sample a batch of transitions $\{s,a,s^{\prime},w\}$ from $\mathcal{D}_{E}$ and $\mathcal{D}_{A}$ Train the Q-function $Q_{\psi}$ using the objective from Eq.8 (CIQL-E) or Eq.9 (CIQL-A): $\psi_{t+1}\leftarrow\psi_{t}+l_{\psi}\cdot\nabla_{\psi}L(Q_{\psi})$ Improve the policy $\pi_{\varphi}$ using the objective from Eq.10 (SAC Actor Update):    $\varphi_{t+1}\leftarrow\varphi_{t}-l_{\varphi}\cdot\nabla_{\varphi}J(\pi_{\varphi})$ Policy $\pi$ interacts with $\mathcal{M}$ and add the new transitions to $\mathcal{D}_{A}$
@@
III-C Dynamics Evaluation
@@
(a) Noise Angle Setting
(b) Initialized
(c) Filtered
(d) Evaluated
@@
Figure 2: Evaluation of Trajectories. $\tau_{1},\tau_{2},\tau_{3}$ are complete trajectories of the approaching stage. $\theta$ is the approach angle between vectors ${\tilde{s}_{t}o}$ and ${\tilde{s}_{t}\tilde{s}_{t+1}}$, and $\theta_{n}$ is the noise angle. (a) and (b) use 15 better trajectories and 15 worse trajectories, and (c) uses 30 better trajectories.
Confidence Score Setting: In the reaching stage, the primary goal is for the robot to approach the target directly. This intention can be represented by the environment state reward function, as shown in Fig.2(a). The blue gradient circle is the operator’s ideal reward function (true intent), where the inner portion is high reward and the outer portion is low reward. To generate this reward function, the useful data is defined as the yellow segments in the trajectory, while the noisy data is defined as the blue segments. Our method is to utilize the approach angle $\theta$ as a classification feature, which is defined as:
$\theta=\arccos\frac{{\tilde{s}_{t}o}\cdot{\tilde{s}_{t}\tilde{s}_{t+1}}}{|{\tilde{s}_{t}o}||{\tilde{s}_{t}\tilde{s}_{t+1}}|}$ (11)
where the value domain of $\theta$ is $[0,\pi]$. $\tilde{s}_{0}$ and $o$ are the initial positions of gripper and target, respectively. The data $\{\tilde{s}_{t},\tilde{s}_{t+1},o\}$ are some features in the transition $\{s_{t},a_{t},s_{t+1}\}$. For a given transition, it has a consistent value $\theta$ that can be used to calculate its confidence score $w$. Since the robot environmental dynamics are deterministic, $w(s_{t},a_{t},s_{t+1})$ is equivalent to $w(s_{t},a_{t})$. Therefore, we can categorize the data into three components based on $\theta$: optimal, non-optimal and noise.
$w(\theta)=\left\{\begin{array}[]{lll}1&for\quad\theta=0,&opt\\ (0,1)&for\quad 0<\theta\leq\theta_{n},&non\ opt\\ 0&for\quad\theta_{n}<\theta\leq\pi,&noise\end{array}\right.$ (12)
where the noise angle $\theta_{n}$ is used to determine whether the data is noisy. Transitions that satisfy $\theta\leq\theta_{n}$ are considered useful, while others are considered noise. The noise-filtered trajectories become well-organized as shown in Fig.2(c). In the useful datasets, it can be further subdivided into optimal and non-optimal data. We find a simple function for $w$ that significantly improves the performance of the original algorithm:
$w(\theta)=\rm Sigmoid((\theta/\theta_{n}-0.5)*\mu)$ (13)
where the ratio $\mu$ satisfies $\rm Sigmoid(0.5*\mu)<\epsilon$, and $\epsilon$ is the limiting distance from the actual score to the confidence bound. The confidence evaluation of the noise-filtered trajectories are shown in Fig.2(d). In the operating stage, it is important to emphasize that it consists of few button-pressing data points. We chose to keep all the data from this stage in the demonstrations where the task can be completed successfully. Their confidence scores are set in the [1, 2] interval to emphasize that stage. Note that the confidence function $w$ relies on dynamics information, eliminating the need for other inputs such as the ground truth rewards or active human supervision.
@@ IV EXPERIMENTS
Figure 3: Teleoperation Demonstration System. Better datasets horizons: 100-150; Worse datasets horizons: 200-400; Failed datasets horizons: 500.
@@ IV-A Experiments Setting
We develop a robotic teleoperation system that leverages the Omega.3 haptic device’s input signaling to gather human demonstrations. This system integrates smoothly with Robosuite, which equips the IIWA robotic arm with an OSC-POSE controller. Control inputs consist of position vectors $(x,y,z)$, angular velocities $(\theta_{p},\theta_{y},\theta_{r})$, and grasping signals $(g)$. We collect three human datasets: Better, Worse, and Failed, each containing 30 trajectories with varying horizons at a 20 Hz control frequency and a horizon limit of 500, as shown in Fig.3.
Training hyperparameters are set to a learning rate of 5e-6 for both actors and critics, and a total of 500K training steps. Training is conducted with a horizon of 200, using the moving average success rate over 10 trajectories to select the best model. Evaluation is based on the average success rate across 100 trajectories from 10 random seeds, using a fixed horizon of 500.
Figure 4: Effect of Noise Angle. IQ-Learn(filter) is learning on filtered noisy data, and it becomes IQ-Learn when $\theta_{n}$ equal to 180°.